Solve 2 cos(θ)=cot(θ) on [0,2π].
Answer by stanbon(75887)

Solve 2 cos(θ)=cot(θ) on [0,2π].
-----
2*cos(t) = cos(t)/sin(t)
----------
2 = 1/sin(t)
sin(t) = 1/2
t = pi/6 or t = (5/6)pi
==========================
Cheers,
Stan H.
--------------------------